OpenStudy (anonymous):

Evaluate the expression for x= 3 3/4 8x + 4

OpenStudy (shamil98):

plug in the value of x that is given into the expression.

OpenStudy (shamil98):

x = 3 3/4 = 15/4 x = 15/4 8x + 4 8(15/4) + 4 = ?

OpenStudy (anonymous):

I think it's 10 but I'm not sure >.>

OpenStudy (shamil98):

It is not ten.

OpenStudy (anonymous):

124?

OpenStudy (shamil98):

\[\huge (8 * \frac{ 15 }{ 4 }) + 4 = ?\] \[\huge ( \cancel 8 2*\frac{ 15 }{ \cancel 4 1 }) + 4 = (2 * 15) + 4 = ?? \]

OpenStudy (anonymous):

34?

OpenStudy (shamil98):

yes

OpenStudy (anonymous):

Thank you